The area bounded by curve y^2 = 4x and the line y = x is:
A. 2/3 sq. units
B. 3/2 sq. units
C. 16/3 sq. units
D. 3/8 sq. units
 

Grade:12th pass

1 Answers

Arun
4 years ago
on solving x = 0 , y = 0 and x= 4, y = 4
 
area = integration of ( y^2/4 – y)  from 0 to 4
 = (y^3/12 – y^2/2) from 0 to 4
 
64/12 – 16/2 = 32/12 = 8/3
 
Hopr it helps